# STKDRec: Spatial-Temporal Knowledge Distillation for Takeaway Recommendation

### v1

@@ Conclusion
In this paper, we propose a novel spatial-temporal knowledge distillation model for takeaway recommendation, termed STKDRec. STKDRec involves two stages: pre-training and STKD stage. During the pre-training stage, we pre-train an STKG encoder to extract rich spatial-temporal knowledge from STKG. During the STKD stage, we apply a ST-Transformer to learn fine-grained spatial region and spatial distance information from the spatial-enhanced sequence perspective. STKD is utilized to integrate heterogeneous spatial-temporal knowledge between STKG and sequence data, reducing training overhead. Experimental results on three real datasets show that the performance of STKDRec is significantly better than the state-of-the-art baselines methods.

### v2

@@ Conclusion
In this paper, we propose a novel spatial-temporal knowledge distillation model for takeaway recommendation, termed STKDRec. The model involves two stages: pre-training and STKD. During the pre-training stage, we train an STKG encoder to extract rich spatial-temporal knowledge from STKG. During the STKD stage, we apply an ST-Transformer to capture dynamic user preferences on fine-grained spatial region and spatial distance information from a sequential perspective. The STKD strategy is utilized to integrate heterogeneous spatial-temporal knowledge from both STKG and sequence data while reducing computational overhead. Experimental results on three real datasets show that the performance of STKDRec is significantly better than the state-of-the-art baselines methods.